### v1

Abstract
We study the $S$-wave $D\bar{D}^{*}(I=0)$ scattering at four different pion masses $m_{\pi}$ ranging from 250 MeV to 417 MeV from $N_{f}=2$ lattice QCD. A bound state near the $D\bar{D}^{*}$ threshold likely exists even after considering the possible left-hand cut effect due to the one pion exchange interaction. At $m_{\pi}\approx 417$ MeV where the effective range expansion is valid, the compositeness of the bound state is $X\approx 1$ and indicates a predominant $D\bar{D}^{*}$ component. This state may correspond to $X(3872)$. On the other hand, our results of the finite volume energies also hint at the existence of a $1^{++}$ resonance below 4.0 GeV with a width around 60 MeV.
Introduction.— Ever since its discovery in 2003 Choi et al. (2003), $X(3872)$ (aka $\chi_{c1}(3872)$ Workman et al. (2022)) has still been a hot topic in experimental and theoretical studies. Its quantum number is set to be $I^{G}J^{PC}=0^{+}1^{++}$ Aaij et al. (2013). $X(3872)$ has a very tiny width of $\sim 1~{}\mathrm{MeV}$ and a mass almost at the $D^{0}\bar{D}^{0*}$ threshold. Apart from its large decay rate to $D^{0}\bar{D}^{0*}$, $X(3872)$ also decays into $J/\psi\omega$ (I=0) and $J/\psi\rho^{0}$ (I=1) with comparable branching fractions Workman et al. (2022). The exotic properties of $X(3872)$ have spurred intensive and extensive phenomenological investigations. $X(3872)$ is unlikely a pure conventional $2P$ charmonium whose mass is expected to be above 3.9 GeV Godfrey and Isgur (1985); Barnes et al. (2005), while its mass and decays suggest it to be a loosely bound $D\bar{D}$ molecule Tornqvist (2004); Liu et al. (2008, 2009); Li and Zhu (2012); Qiu et al. (2023); Dong et al. (2021) (and possibly a small fraction of the $c\bar{c}$ component Braaten and Kusunoki (2004); Kalashnikova (2005); Barnes and Swanson (2008); Ortega et al. (2010); Li et al. (2009); Baru et al. (2010); Yamaguchi et al. (2020)). It is also interpreted as a compact tetraquark state Maiani et al. (2005); Ebert et al. (2006). See Refs. Brambilla et al. (2011); Chen et al. (2016); Esposito et al. (2017); Guo et al. (2018); Brambilla et al. (2020); Liu et al. (2019) for reviews.
There are also several lattice QCD studies considering the $c\bar{c}-D\bar{D}^{(*)}$ coupled channel effects Bali et al. (2011); Prelovsek and Leskovec (2013), where a strong coupling between $c\bar{c}$ and $D\bar{D}^{(*)}$ is observed in the axial vector channel ($J^{PC}=1^{++}$). Especially the study in Ref. Prelovsek and Leskovec (2013) finds that this coupling in the axial vector channel results in a bound state below the $D\bar{D}^{*}$ threshold, which can be related to $X(3872)$. However, the study in Ref. Prelovsek and Leskovec (2013) does not pay enough attention to the role played by the possible $\chi_{c1}(2P)$ charmonium in the interpretation of energy levels above the $D\bar{D}^{*}$ threshold. On the hadron level, the interaction between $D$ and $\bar{D}^{*}$ should be mediated by light hadrons and, therefore, sensitive to light quark masses. In this work, we will investigate the quark mass dependence of the $D\bar{D}^{*}$ interaction, which may shed light on the essence of the related binding mechanism. We will extract the scattering parameters, such as the scattering lengths and the effective ranges at different pion masses. The quark mass dependence can provide more information to the existing effective field theories describing the $D\bar{D}^{*}$ interaction AlFiky et al. (2006); Albaladejo et al. (2013); Baru et al. (2013); Wang and Wang (2013); Guo et al. (2018). In lattice QCD, the scattering parameters are usually inferred from the finite volume energy levels of the $D\bar{D}^{*}$ system through Lüscher’s finite volume method Lüscher (1986, 1991); Luscher (1991). According to the quantum numbers of $X(3872)$, we will focus on the $S$-wave $D\bar{D}^{*}$ scattering in the $I=0$ channel. This is numerically challenging because light quark annihilation diagrams contribute to the related correlation functions. For this purpose, we adopt the distillation method Peardon et al. (2009) that provides a sophisticated treatment of both the all-to-all quark propagators and a smearing scheme for quark fields.
Table 1: Parameters of the gauge ensembles and the energies of $D$ and $D^{*}$ with $q=0,1$. The non-interacting $D\bar{D}^{*}$ energies $E_{D\bar{D}^{*}}^{q=0,1}$ are also shown.
ens. $m_{\pi}$ $a_{t}^{-1}$ $N_{\mathrm{cfg}}$ $N_{V}^{(\mathrm{l})}$ $N_{V}^{(\mathrm{c})}$ $m_{\chi_{c1}}$ $m_{D}$ $m_{D^{*}}$ $E_{D\bar{D}^{*}}^{q=0}$ $E_{D}^{q=1}$ $E_{D^{*}}^{q=1}$ $E_{D\bar{D}^{*}}^{q=1}$
(MeV) (MeV) (MeV) (MeV) (MeV) (MeV) (MeV) (MeV)
M245 250(3) $7.276$ 401 70 120 3489(3) 1873(1) 1985(2) 3858(3) 1958(2) 2064(4) 4022(5)
M305 307(2) $7.187$ 401 70 120 3496(2) 1881(1) 1990(2) 3871(3) 1962(2) 2070(2) 4032(4)
M360 362(1) $7.187$ 401 70 120 3502(2) 1884(1) 2003(2) 3888(2) 1970(1) 2084(3) 4054(4)
M415 417(1) $7.219$ 401 70 120 3509(2) 1896(1) 2017(1) 3913(2) 1978(1) 2094(2) 4072(3)
Numerical details.— We generate $N_{f}=2$ gauge configurations at four pion masses ranging from 250 MeV to 417 MeV on $16^{3}\times 128$ anisotropic lattices with the anisotropy parameter $\xi=a_{s}/a_{t}\approx 5$, where $a_{s}$ and $a_{t}$ are the spatial and the temporal lattice spacings, respectively. We adopt the tadpole improved gauge action Morningstar and Peardon (1997); Chen et al. (2006) for gluons and the tadpole improved anisotropic clover fermion action Zhang and Liu (2001); Su et al. (2006); Meng et al. (2009) for light sea quarks and the valence charm quark. The spatial lattice spacing $a_{s}$ is set to be $a_{s}=0.136(2)~{}\mathrm{fm}$ through the Wilson flow method Lüscher (2010); Borsányi et al. (2012) at the lowest pion mass (the $w_{0}/a_{s}$ is also measured at other pion masses and a mild quark mass dependence is observed with the largest relative deviation being roughly 1-2%). Then we use the measured $\xi$ from the pion dispersion relation to determine $a_{t}$ at each pion mass. The charm quark mass parameter is tuned to give $(m_{\eta_{c}}+3m_{J/\psi})/4=3069$ MeV. Throughout this study, the correlation functions are calculated using the distillation method Peardon et al. (2009). The perambulators for $u,d$ quarks are calculated using $N_{V}^{(\mathrm{l})}=70$ eigenvectors of the gauge covariant Laplacian operators on the lattice, and the perambulators for charm quark are calculated using $N_{V}^{(\mathrm{c})}=120$. The parameters for the gauge ensembles are listed in Table 1.
With respect to the formation of $X(3872)$, we consider the possible coupled channel effects of $\chi_{c1}$ states and the $S$-wave $D\bar{D}^{*}$. In principle, we should consider also the coupling by the $J/\psi\omega$ system, which, however, is observed to almost decouple and is neglected (see Ref. Prelovsek and Leskovec (2013) and also the Supplemental Materials sup ). The interpolating field operators for these states are built in terms of the smeared $u,d,c$ quark fields. The $c\bar{c}$ operators in this Letter takes the spatially extended $c\bar{c}$ version
$\mathcal{O}_{c\bar{c}}^{r}(t)=\frac{1}{N_{r}}\sum\limits_{|\mathbf{y-x}|=r}\bar{c}(\mathbf{x},t)\gamma_{5}\gamma_{i}K_{U}(\mathbf{x},\mathbf{y};t)c(\mathbf{y},t),$ (1)
[…]
For $D\bar{D}^{*}$, we only consider meson-meson operators and ignore the tetraquark-like $qc\bar{c}\bar{q}$ diquark-antidiquark structures. Consider the flavor structure of $D\bar{D}^{*}$ states of the quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$, the $D\bar{D}^{*}$ operators take the form
$\mathcal{O}_{D\bar{D}^{*}}\sim\left[(\bar{u}\Gamma_{D}c)(\bar{c}\Gamma_{D^{*}}u)-(\bar{u}\Gamma_{D^{*}}c)(\bar{c}\Gamma_{D}u)\right]+(u\to d),$ (2)
with $(\Gamma_{D},\Gamma_{D^{*}})=(\gamma_{5},\gamma_{i})$ are used for operators $(\mathcal{O}_{D},\mathcal{O}_{D^{*}})\equiv(\bar{u}(\bar{d})\Gamma_{D}c,\bar{u}(\bar{d})\Gamma_{D^{*}}c)$.
The $S$-wave operator can be obtained by summing over all the possible directions of the relative momentum, namely, $\mathcal{O}_{AB}^{q}\sim\sum\limits_{R\in O}\mathcal{O}_{A}(R\circ\mathbf{q})\mathcal{O}_{B}(-R\circ\mathbf{q})$, where $R$ runs over all the elements of the octahedral group $O$, and $q=|\mathbf{q}|$ is the magnitude of the relative momentum (in units of $2\pi/L$ with $L$ being the spatial size of the lattice). The operators $\mathcal{O}_{4,6}\equiv\mathcal{O}_{D\bar{D}^{*}}^{q=0,1}$ are involved in the practical calculation. Another $\mathcal{O}_{D\bar{D}^{*}}^{q=0}$ operator (labelled as $\mathcal{O}_{5}$) uses the combination $(\Gamma_{D},\Gamma_{D^{*}})=(\gamma_{4}\gamma_{5},\gamma_{4}\gamma_{i})$. This build-up is similar to that in Ref. Prelovsek and Leskovec (2013).
Figure 1: Energy levels $E_{n}$ obtained from GEVP analyses at different $m_{\pi}$. In each panel, the dashed lines show the non-interacting $D\bar{D}^{*}$ energies $E_{D\bar{D}^{*}}^{q=0}$ (green) and $E_{D\bar{D}^{*}}^{q=1}$ (blue) in Table 1, while the boxes stand for the finite volume energy levels $E_{n}$ with their heights indicating $1\sigma$ statistical uncertainties. In the left column of each panel are the lowest two $E_{n}$ when only $c\bar{c}$ operators are involved in the GEVP analysis, and the right column includes the lowest four energy levels $E_{n}$ from the three $c\bar{c}$ operators ($\mathcal{O}_{1,2,3}$) and the three $D\bar{D}^{*}$ operators ($\mathcal{O}_{4,5,6}$).
The six operators $\{\mathcal{O}_{\alpha},\alpha=1,2,\ldots,6\}$ are used to calculate the correlation matrix $C_{\alpha\beta}(t)=\langle 0|\mathcal{O}_{\alpha}(t)\mathcal{O}^{\dagger}_{\beta}(0)|0\rangle$. The practical calculation considers all the quark diagrams, except for those involving the charm quark annihilation. By solving the generalized eigenvalue problem (GEVP), $C_{\alpha\beta}(t)v_{\beta}^{(n)}(t,t_{0})=\lambda^{(n)}(t,t_{0})C_{\alpha\beta}(t_{0})v_{\beta}^{(n)}(t,t_{0})$ at given $t$ and $t_{0}$, we obtain the optimized correlation function $C^{(n)}(t)=v^{(n)}_{\alpha}v^{(n)}_{\beta}C_{\alpha\beta}(t)$ which is contributed mainly by the $n$-th state. We then use the ratio method Aoki et al. (2007); Bulava et al. (2024a, b) to extract the finite volume energies $E_{2,3,4}$: we perform one-exponential fit to the ratio function
$R_{n}(t)=\frac{C^{(n)}(t)}{C_{D}(t,q)C_{D^{*}}(t,q)}\approx Ae^{-\Delta_{n}t},$ (3)
to extract the energy difference $\Delta_{n}=E_{n}-E_{D\bar{D}^{*}}^{q}$, where $C_{D^{(*)}}(t,q)$ is the correlation function of $D^{(*)}$ with a momentum $q=0,1$. In practice, we use $q=0$ for $E_{2,3}$ and $q=1$ for $E_{4}$. We also perform a two-exponential fit to $C^{(n)}(t)$ in a proper time window for a self-consistent check. Although consistent, the energies $E_{2,3,4}$ from the ratio method have smaller errors than those from the two-exponential fit and are therefore used as our final results (See  sup for more details). All the statistical errors are obtained through the jackknife method.
When only the $c\bar{c}$ operators ($\mathcal{O}_{1,2,3}$) are included in the GEVP analysis, for each $m_{\pi}$ we obtain reliably two energy levels, which are shown as yellow and red boxes on the left hand side of each column in Fig. 1, with their heights indicating the statistical errors. The lowest level $E_{1}^{(0)}$ is roughly 3.5 GeV and corresponds necessarily to the conventional charmonium $\chi_{c1}$, while the second level $E_{2}^{(0)}$ is higher than 3.9 GeV and seemingly coincides with the quark model expectation of the $\chi_{c1}(2P)$ mass Godfrey and Isgur (1985); Barnes et al. (2005); Ebert et al. (2003); Li and Chao (2009); Deng et al. (2017); Wang et al. (2019). It should be noted that this energy level does not necessarily correspond to an eigenstate of the lattice Hamiltonian. When the $\mathcal{O}_{D\bar{D}^{*}}$ operators $\mathcal{O}_{4,5,6}$ are added, more energy levels are obtained. The lowest four energy levels (labeled $E_{1},E_{2},E_{3},E_{4}$ from bottom up) are illustrated with colored boxes on the right of each panel in Fig. 1. No sizable change is observed for $E_{1}$ when compared with the lowest level obtained by $c\bar{c}$ operators. The second lowest level $E_{2}$ (green box) appears close to but right below the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=0}$ (green dashed line), and $E_{4}$ lies right above the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=1}$ (blue dashed line), see Table 1 for details. It is interesting to see that $E_{3}$ resides between but far from $E_{D\bar{D}^{*}}^{q=0}$ and $E_{D\bar{D}^{*}}^{q=1}$.
Figure 2: Relative couplings $Z_{\alpha}^{(n)}=\langle 0|\mathcal{O}_{\alpha}|n\rangle$ at $m_{\pi}=417~{}\mathrm{MeV}$. For each operator $\mathcal{O}_{\alpha}$, $Z_{\alpha}^{(n)}$ is normalized by the largest value in $\{Z_{\alpha}^{(n)},n=1,2,\ldots,6\}$. For each state $|n\rangle$, $\{Z_{\alpha}^{(n)},\alpha=1,2,\ldots,6\}$ (in the same color) signal the relative importance of $c\bar{c}$ and $D\bar{D}^{*}$ components.
The energy levels ($E_{1,2,3,4}$) are taken to be the eigenvalues of the lattice Hamiltonian relevant to $0^{+}1^{++}$ charmonium-like systems. The relative couplings of $\mathcal{O}_{\alpha}$ to different states $Z_{\alpha}^{(n)}=|\langle 0|\mathcal{O}_{\alpha}|n\rangle|$ (normalized by the largest value in each panel) are illustrated in Fig. 2. Obviously, the $c\bar{c}$ operators $\mathcal{O}_{1,2,3}$ couple most to the $E_{1}$ state and also have substantial overlaps to the $E_{2}$ and $E_{3}$ states. In contrast, $\mathcal{O}_{4,5}$ ($D\bar{D}^{*}$ operators with $q=0$) couple mainly to $E_{2}$, and also overlap substantially to $E_{1}$ and $E_{3}$ states. $\mathcal{O}_{6}$ ($\mathcal{O}_{D\bar{D}^{*}}(q=1)$ operator) couples predominantly to $E_{4}$ and has a sizable coupling to the $E_{3}$ state. Obviously, the $E_{1}$ state of an energy around $3.5$ GeV is unambiguously assigned to be the $\chi_{c1}(1P)$ state. On the other hand, qualitatively in the language of the state superposition, the $E_{2}$ and $E_{3}$ states are due to the mixing of $\chi_{c1}(2P)$ and non-interacting $D\bar{D}^{*}$ state of $q=0$, which pulls down $E_{2}$ below the $D\bar{D}^{*}$ threshold (apart from the possible attractive interaction of $D\bar{D}^{*}$ due to meson exchanges) and pushes $E_{3}$ upwards, as exactly shown in Fig. 1. It is expected that $E_{4}$ be slightly above the non-interacting energy $E_{D\bar{D}^{*}}^{q=1}$ due to the mixing from $\chi_{c1}(2P)$.
Strictly speaking, $E_{2},E_{3},E_{4}$ are the energy levels of the $D\bar{D}^{*}$ scattering states in a finite box, whose physical implication can be inferred from the corresponding $S$-wave $D\bar{D}^{*}$ scattering amplitude that are conventionally derived through Lüscher’s formula Lüscher (1986, 1991); Luscher (1991). For the single channel $D\bar{D}^{*}$ scattering, one defines for each $E_{n}$ a scattering momentum $p_{n}$ through the relation $E_{n}(p_{n})=\sqrt{m_{D}^{2}+p_{n}^{2}}+\sqrt{m_{D^{*}}^{2}+p_{n}^{2}}$. Then the $S$-wave $D\bar{D}^{*}$ scattering phase shift $\delta_{0}(p_{n})$ reads
$p\cot\delta_{0}(p)=\frac{2}{\sqrt{\pi}L}\mathcal{Z}_{00}(1;q^{2}),~{}~{}~{}q^{2}\equiv\left(\frac{L}{2\pi}\right)^{2}p^{2},$ (4)
Table 2 collects the results of $(p^{2},p\cot\delta_{0}(p))$ of $E_{2,3,4}$ at all the four values of $m_{\pi}$, where the statistical errors are obtained by jackknife analyses.
However, there are subtleties in the derivation of $p\cot\delta_{0}$ for $E_{2}$. The dynamics of the $D\bar{D}^{*}$ scattering may includes the interaction from one pion exchange (OPE), which introduces nonanalyticity to $p\cot\delta_{0}(p)$ when $p^{2}$ is below the left-hand cut (lhc) point $(p_{\mathrm{lhc}}^{1\pi})^{2}=((m_{D^{*}}-m_{D})^{2}-m_{\pi}^{2})/4$ Du et al. (2023); Meng et al. (2023); Raposo and Hansen (2023).
Figure 3: Phase shifts of S-wave $J^{PC}=1^{++}$ $D\bar{D}^{*}$ system and effective range expansion fitting. The red data points illustrate phase shifts $p\cot\delta_{0}(q)$ obtained from ($E_{2,3}$), while the green bands show the related effective ranges expansion. The background dashed lines present the Riemann-Zeta function trajectory.
Table 2: The $D\bar{D}^{*}$ scattering parameters. The scattering momentum $p^{2}$, the phase shift $p\cot\delta_{0}(p)$ (and $\delta_{0}$) at $E_{2,3,4}$, as well as the lhc point $(p_{\mathrm{lhc}}^{1\pi})^{2}\approx\frac{1}{4}[(m_{D}-m_{D^{*}})^{2}-m_{\pi}^{2}]$, are given explicitly for all the four values of $m_{\pi}$. The energy shifts $\Delta_{n}$ (derived from Eq. (3)) of $E_{n}$ from the nearby non-interacting energy of $D\bar{D}^{*}$ are also presented. Also collected are the derived scattering length $a_{0}$, the effective range $r_{0}$, and the estimated binding energy $E_{B}$ for the bound state pole. The asterisks in brackets ‘(*)’ indicate that the corresponding values may be problematic when OPE lhc is considered.
$m_{\pi}$(MeV) 250(3) 307(2) 362(1) 417(1)
$E_{4}$ $\Delta_{4}$(MeV) $=E_{4}-E_{D\bar{D}^{*}}^{q=1}$ 9.1(1.3) 8.9(1.2) 5.3(1.3) 12.8(1.3)
$p^{2}(\mathrm{GeV}^{2})$ 0.339(8) 0.335(6) 0.340(6) 0.342(4)
$p\cot\delta_{0}(p)$(GeV) -2.02(66) -2.35(65) -2.76(89) -1.79(28)
$\delta_{0}$ $(163.9_{-7.3}^{+4.0})^{\circ}$ $(166.1_{-5.1}^{+3.0})^{\circ}$ $(168.1_{-5.5}^{+2.9})^{\circ}$ $(161.9_{-3.2}^{+7.4})^{\circ}$
$E_{3}$ $\Delta_{3}$(MeV)$=E_{3}-E_{D\bar{D}^{*}}^{q=0}$ 70(3) 63(3) 80(3) 80(3)
$p^{2}(\mathrm{GeV}^{2})$ 0.135(5) 0.122(6) 0.158(6) 0.158(6)
$p\cot\delta_{0}(p)$(GeV) -0.054(19) -0.097(19) 0.012(22) 0.026(24)
$\delta_{0}$ $(98.4_{-3.6}^{+3.3})^{\circ}$ $(105.4_{-3.1}^{+3.0})^{\circ}$ $(88.2_{-3.3}^{+3.3})^{\circ}$ $(86.2_{-4.1}^{+4.0})^{\circ}$
$E_{2}$ $\Delta_{2}$(MeV)$=E_{2}-E_{D\bar{D}^{*}}^{q=0}$26.1(9) -25.4(11) -19.0(7) -18.6(8)
$p^{2}(\mathrm{GeV}^{2})$0.050(2) -0.049(2) -0.037(1) -0.036(1)
$p\cot\delta_{0}(p)$(GeV) -0.154(9)(*) -0.146(10)(*) -0.063(11)(*) -0.066(13)
$(p_{\mathrm{lhc}}^{1\pi})^{2}(\mathrm{GeV}^{2})$0.0135(4) -0.0210(4) -0.0292(3) -0.0400(3)
$a_{0}$ (fm) -1.55(10)(*) -1.50(12)(*) -4.03(91)(*) $\mathbf{-4.0(1.0)}$
$r_{0}$ (fm) 0.211(30)(*) 0.113(34)(*) 0.153(34)(*) $\mathbf{0.187(38)}$
$E_{B}$ (MeV) $-9.7^{+2.1}_{-2.2}$ (*) $-9.7^{+1.9}_{-2.0}$ (*) $-1.3^{+0.6}_{-0.8}$ (*) $\mathbf{-1.3^{+0.8}_{-1.0}}$
Bound state and the possible lhc issue.— The values of $(p_{\mathrm{lhc}}^{1\pi})^{2}$ at different $m_{\pi}$’s are also shown in Table 2. It is seen that $p^{2}(E_{2})>(p_{\mathrm{lhc}}^{1\pi})^{2}$ is satisfied only for $m_{\pi}=417(1)$ MeV. So we start the discussion of the near-threshold $D\bar{D}^{*}$ scattering amplitude at this $m_{\pi}$ following the strategy in Ref. Prelovsek and Leskovec (2013); Mohler et al. (2013), since the $p\cot\delta_{0}(p)$ derived at $E_{2}$ and $E_{3}$ from Eq. (4) is valid. The effective range expansion (ERE), $p\cot\delta_{0}(p)=\frac{1}{a_{0}}+\frac{1}{2}r_{0}p^{2}$, at $E_{2}$ and $E_{3}$ gives two equations with two unknowns, namely, the $S$-wave scattering length $a_{0}$ and the effect range $r_{0}$. The solutions are
$a_{0}=-4.0(1.0)~{}\mathrm{fm},~{}~{}~{}r_{0}=0.187(38)~{}\mathrm{fm},$ (5)
where the large negative $a_{0}$ signals the existence of a bound state. Taking these values as the approximation of those in the infinite volume, we can solve the pole equation of the scattering amplitude $\mathcal{T}\propto(p\cot\delta_{0}(p)-ip)^{-1}$ to obtain the pole momentum $p_{B}=i\kappa_{B}$ with $\kappa_{B}=49.8_{-18.2}^{+17.0}$ MeV. Consequently, we get the binding energy $E_{B}$,
$E_{B}=E_{D\bar{D}^{*}}(p_{B})-(m_{D}+m_{D^{*}})=-1.3_{-1.0}^{+0.8}~{}\mathrm{MeV}.$ (6)
The Weinberg compositeness criterion states that a near-threshold bound state can be viewed as an admixture of a compact state and a two-hadron channel. The generalized Weinberg relation in Ref. Li et al. (2022), which is more proper for the $D\bar{D}^{*}$ single-channel problem here, implies that the compositeness $X$ computed with ERE should be $X\approx 1$ up to ${O}(p^{2})$ according to the small positive $r_{0}$. Therefore, this bound state is predominantly a $D\bar{D}^{*}$ bound state.
If the same procedure is applied to cases at the other three lighter $m_{\pi}$’s, then similar results can be obtained, as shown in Table 2. The “binding energy” $E_{B}=-9.7_{-2.2}^{+2.1}$ at $m_{\pi}=250(3)$ MeV is also consistent with the value $11(7)$ MeV obtained in Ref. Prelovsek and Leskovec (2013) at $m_{\pi}\approx 266$ MeV. However, the $p^{2}$ of $E_{2}$ for these $m_{\pi}$’s are lower than the corresponding lhc point $(p_{\mathrm{lhc}}^{1\pi})^{2}$, such that the $p\cot\delta_{0}(p)$ at $E_{2}$ derived by Eq. (4) is questionable. In Refs. Du et al. (2023); Meng et al. (2023), case studies are performed to the OPE lhc effect on the $DD^{*}$ scattering relevant to $T_{cc}^{+}(3875)$ using the lattice data obtained in Ref. Padmanath and Prelovsek (2022). After considering the OPE, these effective theory studies show that, except for the singular behavior close to $(p_{\mathrm{lhc}}^{1\pi})^{2}$ that $p\cot\delta_{0}(p)$ bends up abruptly when $p^{2}$ approaching $(p_{\mathrm{lhc}}^{1\pi})^{2}$ from above, the overall behavior with respect to $p^{2}$ is similar to that of ERE including the values of $p^{2}$ below $(p_{\mathrm{lhc}}^{1\pi})^{2}$. Figure 3 shows the phase shifts (red points), as well as ERE of $p\cot\delta_{0}(p)$ (green bands) for the four $m_{\pi}$’s. In each panel, the ERE curve intersects with the bound state curve $ip$ (dashed green line). If the $D\bar{D}^{*}$ scattering also has the similar feature to that of the $DD^{*}(I=0)$ scattering when considering the OPE lhc, then it is possible that $p\cot\delta_{0}(p)$ has also a pole near the lhc point $(p_{\mathrm{lhc}}^{1\pi})^{2}$ (red dashed vertical line) and exhibits a rapid increase there to intersect earlier with the bound state curve $ip$ at $p^{2}>(p_{\mathrm{lhc}}^{1\pi})^{2}$. Therefore, it is very likely that there exists a bound state pole at each $m_{\pi}$ with binding energy $E_{B}$ larger ($|E_{B}|$ smaller) than the value in Table 2.
Possible resonance.— Now we consider the physical implication of the energy level $E_{3}$, which lies between but far from the $D\bar{D}^{*}$ threshold and $E_{D\bar{D}^{*}}^{q=1}$. In Ref. Prelovsek and Leskovec (2013), the $E_{3}$ state is interpreted as the upward shifted $D\bar{D}^{*}$ state with the relative momentum $q=0$ owing to the formation of a bound state below the threshold. The physical consideration behind this is Levinson’s theorem that, when $n_{l}$ bound states are formed in the $l$ partial wave, the phase shift of the $l$-wave phase shift satisfies the relation $\delta_{l}(p=0^{+})=n_{l}\pi$ with the convention $\delta(p=\infty)=0$. However, when there exist $n_{b}$ bare states between the threshold and a specific momentum $p_{\mathrm{max}}$, one should use the generalized Levinson’s theorem $\delta_{l}(0^{+})-\delta_{l}(p_{\mathrm{max}})=(n_{l}-n_{b})\pi$ Vidal and LeTourneux (1992); Li and Wu (2022). For the case at hand with $n_{0}=1$, if the bare $\chi_{c1}(2P)$ state exists theoretically, then the phase shift $\delta_{0}(p)$ will undergo an evolution that it starts from $\pi$ at $p=0^{+}$, falls down when the energy increases, and then return to $\pi$ when the energy passes the $\chi_{c1}(2P)$ mass, as is exactly the case (see the value of the phase shift $\delta_{0}$ at $E_{4}$ in Table 2). It is stressed that the position of $E_{4}$ relative to $E_{D\bar{D}^{*}}^{q=1}$ is crucial for this reasoning. By using the ratio method in Eq. (3), we obtain $E_{4}$ right above $E_{D\bar{D}^{*}}^{q=1}$. Actually, this phase change also hints the existence of a resonance. The Breit-Wigner ansatz for a resonance $(m_{R},\Gamma_{R})$ gives
$\delta_{0}(E)=\arctan\left(\frac{\Gamma_{R}}{2(m_{R}-E)}\right).$ (7)
Using $\delta_{0}(E_{3})$ and $\delta_{0}(E_{4})$ as inputs, the resonance paramters $(m_{R},\Gamma_{R})$ at each $m_{\pi}$ are estimated and are collected in Table 3. Figure 4 illustrates the evolution of $\delta_{0}(E)$ with respect to the center-of-mass energy $E$ of $D\bar{D}^{*}$, where red points are the lattice data, the blue bands are plots of Eq. (7) with parameters in Table 3, and the dashed lines indicate the possible phase change below the resonant energy $E_{R}$.
Figure 4: Fitting of resonance using the Breit-Wigner ansatz. In each panel of the figure the red data points represent the phase shifts corresponding to $E_{3}$ and $E_{4}$ in Table 2 at different $m_{\pi}$. The blue band illustrated the resonance using parameters $(m_{R},\Gamma_{R})$. The blue dashed lines illustrate the schematic evolution of phase shifts starting from $\delta_{0}(0^{+})\sim 180^{\circ}$ such that $\delta_{0}(0^{+})-\delta_{0}(p_{\mathrm{max}})\to 0$ when the $D\bar{D}^{*}$ energy $E(p_{\mathrm{max}})$ passes the mass of the would-be $\chi_{c1}(2P)$.
Table 3: Resonance parameters $(m_{R},\Gamma_{R})$.
$m_{\pi}$(MeV) 250(3) 307(2) 362(1) 417(1)
$m_{R}$(MeV) 3924(5) 3926(6) 3969(4) 3995(4)
$\Gamma_{R}$(MeV) 63(23) 57(18) 37(13) 57(10)
Phenomenological studies also indicate that when the coupled channel effect of $D\bar{D}^{*}$ and the quark model $\chi_{c1}(2P)$ is considered, apart from a narrow structure near the $D^{0}\bar{D}^{*0}$ threshold, there may exist a wide resonance below 4.0 GeV with the width ranging from 17 to 80 MeV Kalashnikova (2005); Cincioglu et al. (2016); Giacosa et al. (2019); Deng et al. (2023); Wang et al. (2023). These observations are compatible with the results of this study. Experimentally, $X(3940)$ reported by Belle in 2007  Abe et al. (2007); Pakhlov et al. (2008) has the resonance parameters $(m_{X},\Gamma)\sim(3942(9),37_{-17}^{+27})$ MeV, and decays mainly into $D\bar{D}^{*}$ but not $D\bar{D}$, and therefore favors the assignment of a $1^{++}$ resonance. However, it is puzzling that a similar structure has not been observed by other experiments yet.
Summary:— We calculate the $I^{G}J^{PC}=0^{+}1^{++}$ channel $D\bar{D}^{*}$ scattering at four pion masses ranging from 250 MeV to 417 MeV in $N_{f}=2$ lattice QCD. We use the distillation method to calculate the related correlation matrix of $c\bar{c}$ operators and $D\bar{D}^{*}$ operators. At $m_{\pi}\approx 417$ MeV where there are no severe OPE lhc issues, a shallow bound state is observed to reside below the $D\bar{D}^{*}$ by $1.3_{-1.0}^{+0.8}$ MeV, and is a $D\bar{D}^{*}$ component of 96(1)%. When considering the OPE lhc effect, the results at the lighter pion masses also indicate the possible existence of a bound state. In the meantime, the finite volume energy levels above the $D\bar{D}^{*}$ threshold hint the existence of a $1^{++}$ resonance below $4.0$ GeV with a width around $60$ MeV. This is compatible with the expectation of phenomenological studies.
Acknowledgements:— We are grateful to Profs. Qiang Zhao, Feng-Kun Guo, Jia-Jun Wu and Xu Feng for the valuable discussions. This work is supported by the National Natural Science Foundation of China (NNSFC) under Grants No. 11935017, No. 12293060, No. 12293061, No. 12293065, No. 12293063, No. 12075253, No. 12070131001 (CRC 110 by DFG and NNSFC), No. 12175063, No. 12205311, the National Key Research and Development Program of China (No. 2020YFA0406400) and the Strategic Priority Research Program of Chinese Academy of Sciences (No. XDB34030302). The Chroma software system Edwards and Joo (2005) and QUDA library Clark et al. (2010); Babich et al. (2011) are acknowledged. The computations were performed on the HPC clusters at the Institute of High Energy Physics (Beijing) and China Spallation Neutron Source (Dongguan), and the ORISE computing environment.
S1. Conventions and Operators
We take the following procedure to build $D\bar{D}^{*}$ operators with quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$. According to the isospin assignment of $u,d$ quarks,
$\begin{cases}|u\rangle=\left|\frac{1}{2},\frac{1}{2}\right\rangle,\\ |d\rangle=\left|\frac{1}{2},-\frac{1}{2}\right\rangle,\end{cases}\begin{cases}|\bar{d}\rangle=\left|\frac{1}{2},\frac{1}{2}\right\rangle,\\ |\bar{u}\rangle=-\left|\frac{1}{2},-\frac{1}{2}\right\rangle,\end{cases}$ (SI1)
we have the isospin states of $D$ and $D^{*}$,
$\begin{cases}|D^{(*)+}\rangle=|c\bar{d}\rangle=|\frac{1}{2},\frac{1}{2}\rangle,\\ |D^{(*)0}\rangle=|c\bar{u}\rangle=-|\frac{1}{2},-\frac{1}{2}\rangle,\\ |\bar{D}^{(*)0}\rangle=|u\bar{c}\rangle=|\frac{1}{2},\frac{1}{2}\rangle,\\ |D^{(*)-}\rangle=|d\bar{c}\rangle=|\frac{1}{2},-\frac{1}{2}\rangle.\\ \end{cases}$ (SI2)
Thus the $I=0$ combination of charge neutral $D\bar{D}^{*}$ states reads
$|D\bar{D}^{*}\rangle_{I=0}^{Q=0}=\frac{1}{2}\left(|D^{+}\bar{D}^{*-}\rangle+|D^{0}\bar{D}^{*0}\rangle-|\bar{D}^{0}D^{*0}\rangle-|D^{-}D^{*+}\rangle\right).$ (SI3)
Since the charge conjugation transformation ($\mathcal{C})$ of $D$ and $D^{*}$ are conventionally defined as $\mathcal{C}|D\rangle=+|\bar{D}\rangle$, $\mathcal{C}|D^{*}\rangle=-|\bar{D}^{*}\rangle$, one can easily check the $\mathcal{C}$ transformation $\mathcal{C}|D\bar{D}^{*}\rangle_{I=0}^{Q=0}=+|D\bar{D}^{*}\rangle_{I=0}^{Q=0}$. The $\mathcal{P}$-parity $P=+$ requires the $D\bar{D}^{*}$ relative angular momentum $l=0$. Therefore, the $S$-wave $D\bar{D}^{*}$ state with the combination in Eq. (SI3) has the desired quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$. The combination in Eq. (SI3) gives also the flavor structure of the interpolation field operator of $D\bar{D}^{*}$ systems.
A $S$-wave two-particle operator $\mathcal{O}_{AB}(q)$ is easily constructed by the summing over all the relative momenta $\mathbf{p}\equiv 2\pi\mathbf{q}/L$ with the same $q=|\mathbf{q}|$, namely,
$\mathcal{O}_{AB}^{q}=\frac{1}{N_{q}}\sum\limits_{R\in O}\mathcal{O}_{A}(R\circ\mathbf{q})\mathcal{O}_{B}(-R\circ\mathbf{q}).$ (SI4)
where $R$ runs over all the elements of the octahedral group $O$ and $N_{q}$ is the degeneracy of $q$. Given the single-particle operators of $D$ and $\bar{D}^{*}$, $(\mathcal{O}_{D},\mathcal{O}_{D^{*}})\equiv(\bar{u}(\bar{d})\Gamma_{D}c,\bar{u}(\bar{d})\Gamma_{D^{*}}c)$, we combine the relations in Eq. (SI3) and (SI4) to build up the $D\bar{D}^{*}$ operators $\mathcal{O}_{D\bar{D}^{*}}^{q}$ with the quantum numbers $I^{G}J^{PC}=0^{+}1^{++}$. In practice, we use two combinations of $\Gamma_{D}$ and $\Gamma_{D^{*}}$, namely, $(\Gamma_{D},\Gamma_{D^{*}})=(\gamma_{5},\gamma_{i})$ and $(\gamma_{4}\gamma_{5},\gamma_{4}\gamma_{i})$. The operators $\mathcal{O}_{D\bar{D}^{*}}^{q=0,\gamma_{5}(\gamma_{4}\gamma_{5})}$ with the two $(\Gamma_{D},\Gamma_{D^{*}})$ combinations (indicated by the superscripts $\gamma_{5}$ and $\gamma_{4}\gamma_{5}$) are also named by $\mathcal{O}_{4}$ and $\mathcal{O}_{5}$, respectively. The $\mathcal{O}_{D\bar{D}^{*}}^{q=1}$ operator uses the first $(\Gamma_{D},\Gamma_{D^{*}})$ combination and is also named as $\mathcal{O}_{6}$.
The $S$-wave $J/\psi\omega$ operator with $q=0$ is simply the antisymmetric combination of the single particle operators $\mathcal{O}_{J/\psi}^{j}(q=0)$ and $\mathcal{O}_{\omega}^{k}(q=0)$, namely, $\mathcal{O}_{7}\equiv\mathcal{O}_{J/\psi\omega}^{q=0}\to\epsilon_{ijk}\mathcal{O}_{J/\psi}^{j}(q=0)\mathcal{O}_{\omega}^{k}(q=0)$.
For the convenience of discussions, we recite the build-up of $c\bar{c}$ operators presented in the main text. Since the $\chi_{c1}$ mass $m_{\chi_{c1}}\approx 3.51~{}\mathrm{GeV}$ is far from the energy region relevant to $X(3872)$, we need several $c\bar{c}$ operators to distinguish states of interest from $\chi_{c1}$. So we build several spatially extended charmonium $c\bar{c}$ operators
$\mathcal{O}_{c\bar{c}}^{r}(t)=\frac{1}{N_{r}}\sum\limits_{|\mathbf{y-x}|=r}\bar{c}(\mathbf{x},t)\gamma_{5}\gamma_{i}K_{U}(\mathbf{x},\mathbf{y};t)c(\mathbf{y},t),$ (SI5)
where $N_{r}$ is the multiplicity of $\mathbf{r}=\mathbf{y-x}$ with $|\mathbf{r}|=r$, and
[…]
is a Wilson line connecting $(\mathbf{y},t)$ and $(\mathbf{x},t)$. Obviously, $\mathcal{O}_{c\bar{c}}^{r}(t)$ is gauge invariant and has the right quantum number $J^{PC}=1^{++}$ (actually $T_{1}^{++}$ on the lattice) after the summation over $|\mathbf{r}|=r$. In practice, we use three $\mathcal{O}_{c\bar{c}}^{r}(t)$ operators (denoted by $\mathcal{O}_{1,2,3}$) with $r/a_{s}=0,1,2$, respectively.
Figure SI1: Schematic quark contraction diagrams for the calculation of the correlation matrix $C_{\alpha\beta}(t)$. For convenience, the operator set is sorted as $\{\mathcal{O}_{\alpha}|\alpha=1,\cdots,7\}=\{\mathcal{O}_{c\bar{c}}^{r=0},\mathcal{O}_{c\bar{c}}^{r=1},\mathcal{O}_{c\bar{c}}^{r=2},\mathcal{O}_{D\bar{D}}^{\mathbf{q}=0,\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=0,\gamma_{4}\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=1},\mathcal{O}_{J/\psi\omega}^{{q}=0}\}$. All the contributions are considered except for charm quark annihilation diagrams which are expected to be suppressed by the OZI rule.
Figure SI2: Effective mass using different operators subset and solving GEVP, ranging from different $m_{\pi}$. In each column of the panels, the headers indicate the operators involved in GEVP. The colored lines show the effective mass behavior corresponding to their dominant operators, as indicated by the same color. Their fitting values are given in Table.[SI1-SI4].
S2. Determination of finite volume energies
Based on the operator set introduced above,
$\mathcal{S}=\{\mathcal{O}_{\alpha}|\alpha=1,\cdots,7\}=\{\mathcal{O}_{c\bar{c}}^{r=0},\mathcal{O}_{c\bar{c}}^{r=1},\mathcal{O}_{c\bar{c}}^{r=2},\mathcal{O}_{D\bar{D}}^{\mathbf{q}=0,\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=0,\gamma_{4}\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=1},\mathcal{O}_{J/\psi\omega}^{{q}=0}\},$ (SI7)
we calculate the corresponding correlation matrix
$C_{\alpha\beta}(t)=\frac{1}{T}\sum\limits_{\tau}\langle\mathcal{O}_{\alpha}(t+\tau)\mathcal{O}_{\beta}(\tau)\rangle~{}~{}(\alpha,\beta=1,2,\ldots,7),$ (SI8)
where the correlation functions with different source time slices $\tau$ are averaged to increase the statistics. Figure SI1 shows the quark diagrams after Wick’s contraction that are involved in the calculation of $C_{\alpha\beta}(t)$. All the connected quark diagrams and the diagrams including the light quark annihilation effects are taken into account, while the charm quark annihilation effects are neglected owing to the OZI rule.
First, we take the following procedure to check the relevance of individual operators $\mathcal{O}_{\alpha}\in\mathcal{S}$ to the $D\bar{D}^{*}$ scattering we are interested in:
  • We start with the subset $\{\mathcal{O}_{1},\mathcal{O}_{2},\mathcal{O}_{3}\}$ ($c\bar{c}$ operators) of $\mathcal{S}$, whose correlation matrix $\{C_{ij}(t),i,j=1,2,3\}$ is actually a submatrix $\{C_{ij},i,j=1,2,3\}$ of $C_{\alpha\beta}(t)$. We solve the generalized eigenvalue problem (GEVP) to the submatrices of $C_{ij}(t)$,
    $C_{ij}(t_{1})v_{j}^{(n)}(t_{1},t_{0})=\lambda^{(n)}(t_{1},t_{0})C_{ij}(t_{0})v_{j}^{(n)}(t_{1},t_{0}),$ (SI9)
    for given $t_{1},t_{0}$. Thus we obtain the optimized correlation functions $C^{(n)}(t)=v_{i}^{(n)}(t_{1},t_{0})v_{j}^{(n)}(t_{1},t_{0})C_{ij}(t)$. This procedure runs over the cases at the four $m_{\pi}$’s. The effective masses of $C^{(1)}(t)$ and $C^{(2)}(t)$ are illustrated in the rightmost column (column (e)) in each panel of Fig. SI2. When $\mathcal{O}_{7}(J/\psi\omega,q=0)$ is added to the operator subset, the solution to the corresponding GEVP gives an additional energy level close to the $J/\psi\omega$ threshold, as shown in column (c) as black points in each panel of Fig. SI2.
  • Now we perform the GEVP analysis to the operator subset $\{\mathcal{O}_{4},\mathcal{O}_{5},\mathcal{O}_{6},\mathcal{O}_{7}\}$. The effective masses of the four optimized correlation functions are plotted in column (d) of each panel of Fig. SI2, where the effective mass illustrated by black points is almost the same as that in column (c). Other three effective masses have good signal-to-noise ratios but do not show plateaus even in the time range $t/a_{t}\in[25,40]$.
  • If the GEVP analysis is performed to the full operator set $\mathcal{S}$, we obtain seven optimized correlation functions. The effective masses of the lowest six states are plotted in column (a) of each panel of Fig. SI2. One can see that the effective masses of the lowest five states tend to saturate on plateaus. Column (b) in each panel of Fig. SI2 shows the effective masses when the $J/\psi\omega$ operator is excluded from the operator set. In comparison with column (a), the state illustrated by black points disappears but other states are still there and almost do not change.
The observations above indicate that the $J/\psi\omega$ operator ($\mathcal{O}_{7}$) almost does not couple with $c\bar{c}$ and $D\bar{D}^{*}$ operators and are therefore nearly irrelevant to the $D\bar{D}^{*}$ scattering. In contrast, $c\bar{c}$ and $D\bar{D}^{*}$ operators strongly couples with each other and must be considered together. These observations are in agreement with the previous work Ref. Prelovsek and Leskovec (2013).
As for the energy levels $E_{n}$ to be determined, we take the convention that the energy levels reflected by the effective masses in column (b) of Fig. SI2 are ordered as $E_{1},E_{2},\ldots,E_{6}$ from bottom to top. For a specific subset of the full operator set $\mathcal{S}$, the energy level that is adjacent to $E_{n}$ has the same index $n$. We do not include the energy level (indicated by black points) corresponding to the $J/\psi\omega$ to the level ordering.
For each specific operator subset, we first perform two-exponential fits to the optimized correlation functions $C^{(n)}(t)$ introduced above and take the lowest energy to be an estimate of $E_{n}$, which are listed in Tables SI1-SI4 for all the four $m_{\pi}$’s. In the data blocks in each table, the first two rows of the leftmost block are almost the same and are the energy levels $E_{n}$ from $c\bar{c}$ operators with and without the $J/\psi\omega$ operator. The energies $E_{n}$ in first two rows of the middle block are from $D\bar{D}^{*}$ operators with and without the $J/\psi\omega$ operator. The energies $E_{n}$ in the last two rows are from both $c\bar{c}$ and $D\bar{D}^{*}$ operators (with and without the $J/\psi\omega$ operator. The energy values in bold are determined from the ratio method (see below). It is clearly seen that the inclusion or exclusion of the $J/\psi\omega$ operator almost does not affect other energy values. So we exclude the $J/\psi\omega$ operator in the following discussions.
Comparing the energy levels from only $c\bar{c}$ operators, those from only $*D\bar{D}*$ operators, and those from both $c\bar{c}$ and $D\bar{D}^{*}$ operators, the strong correlation between $c\bar{c}$ and $D\bar{D}^{*}$ operators is obviously manifested. Especially, the three energy levels from only $*D\bar{D}*$ operators are far from non-interacting $D\bar{D}^{*}$ energies. This is attributed to the light quark annihilation effects illustrated in Fig. SI1. Therefore, it is necessary to include $c\bar{c}$ operators when the $I=0$ $D\bar{D}^{*}$ scattering is considered. In this sense, the energies $E_{n}$ (the bottom row in Tables SI1-SI4) from both both $c\bar{c}$ and $D\bar{D}^{*}$ operators are taken as good estimates of the eigenvalues of the lattice Hamiltonian relevant to the $I^{G}J^{PC}=0^{+}1^{++}$ charmonium-like system.
Table SI1: Energy levels $E_{n}$ from different subsets of operators for lattice ensemble $m_{\pi}=250~{}\mathrm{MeV}$. In each row of $E_{n}$, the energy values (converted to the physical unit GeV through the fixed $a_{t}^{-1}$ in Table SI5) are obtained from the operators in the non-empty columns. Therefore, the first four rows of $E_{n}$ show the unmixed energies. The mixing effects are very weak and can almost be neglected for $D\bar{D}^{*}(q=0)-D\bar{D}^{*}(q=1)$ (row no.5), $J/\psi\omega-D\bar{D}^{*}$ (row no.6) and $J/\psi\omega-c\bar{c}$ (row no.7) mixings. The strong mixing takes place between $c\bar{c}$ states and $D\bar{D}^{*}$ states, which is manifested by the substantial changes of energies $E_{2}$ and $E_{3}$ after the mixing (row no.8 and row no.9).
$\mathcal{O}_{\alpha}$ $\mathcal{O}_{1,2,3}$ $\mathcal{O}_{4}(D\bar{D}^{*},q=0)$ $\mathcal{O}_{5}(D\bar{D}^{*},q=0)$ $\mathcal{O}_{6}(D\bar{D}^{*},q=1)$ $\mathcal{O}_{7}(J/\psi\omega,q=0)$
$n$ 1 3 5 2 4 6 -
$E_{n}$(GeV) 3.4907(20) 3.9082(47) 4.087(54)
3.4907(20) 3.9082(47) 4.087(54) 3.889(13)
3.5099(65) 4.142(22) 4.215(49)
3.5099(65) 4.142(22) 4.215(49) 3.889(13)
3.4896(19) 3.9282(46) 4.089(55) 3.8324(36) 4.0308(49) 4.577(14) 3.889(13)
3.4896(19) 3.9282(46) 4.089(55) 3.8324(36) 4.0308(49) 4.577(14)
Table SI2: Similar to Tab. SI1 but for $m_{\pi}=307~{}\mathrm{MeV}$.
$\mathcal{O}_{\alpha}$ $\mathcal{O}(r=0,a_{s},2a_{s})$ $\mathcal{O}_{D\bar{D}^{*}}^{1}(q=0)$ $\mathcal{O}_{D\bar{D}^{*}}^{2}(q=0)$ $\mathcal{O}_{D\bar{D}^{*}}^{1}(q=1)$ $J/\psi\omega$
$n$ 1 3 5 2 4 6 -
$E_{n}$(GeV) 3.4947(20) 3.9135(47) 4.178(45)
3.4947(20) 3.925(18) 4.178(45) 3.878(13)
3.5180(67) 4.135(24) 4.133(47)
3.5180(67) 4.135(24) 4.133(47) 3.878(13)
3.4935(19) 3.9343(43) 4.182(45) 3.8461(33) 4.0411(36) 4.619(13) 3.878(13)
3.4935(19) 3.9343(43) 4.182(45) 3.8461(33) 4.0411(36) 4.619(13)
Table SI3: Similar to Tab. SI1 but for $m_{\pi}=362~{}\mathrm{MeV}$.
$\mathcal{O}_{\alpha}$ $\mathcal{O}(r=0,a_{s},2a_{s})$ $\mathcal{O}_{D\bar{D}^{*}}^{1}(q=0)$ $\mathcal{O}_{D\bar{D}^{*}}^{2}(q=0)$ $\mathcal{O}_{D\bar{D}^{*}}^{1}(q=1)$ $J/\psi\omega$
$n$ 1 3 5 2 4 6 -
$E_{n}$(GeV) 3.5021(23) 3.9528(44) 4.242(41)
3.5021(23) 3.9528(44) 4.242(41) 3.923(10)
3.5208(61) 4.131(21) 4.187(37)
3.5208(61) 4.131(21) 4.187(37) 3.923(10)
3.5018(21) 3.9680(40) 4.245(41) 3.8687(26) 4.0589(36) 4.688(11) 3.923(10)
3.5018(21) 3.9680(40) 4.245(41) 3.8687(26) 4.0589(36) 4.688(11)
Table SI4: Similar to Tab. SI1 but for $m_{\pi}=417~{}\mathrm{MeV}$.
$\mathcal{O}_{\alpha}$ $\mathcal{O}(r=0,a_{s},2a_{s})$ $\mathcal{O}_{D\bar{D}^{*}}^{1}(q=0)$ $\mathcal{O}_{D\bar{D}^{*}}^{2}(q=0)$ $\mathcal{O}_{D\bar{D}^{*}}^{1}(q=1)$ $J/\psi\omega$
$n$ 1 3 5 2 4 6 -
$E_{n}$(GeV) 3.5089(21) 3.9778(43) 4.121(73)
3.5089(21) 3.9778(43) 4.121(73) 3.9688(77)
3.5320(55) 4.179(24) 4.348(70)
3.5320(55) 4.179(24) 4.348(70) 3.9688(77)
3.5083(20) 3.9934(41) 4.123(74) 3.8949(24) 4.0850(27) 4.720(11) 3.9688(77)
3.5083(20) 3.9934(41) 4.123(74) 3.8949(24) 4.0850(27) 4.720(11)
Figure SI3: Relative couplings $Z_{\alpha}^{(n)}=\langle 0|\mathcal{O}_{\alpha}|n\rangle$ at various $m_{\pi}$. For each operator $\mathcal{O}_{\alpha}$, $Z_{\alpha}^{(n)}$ is normalized by the largest value in $\{Z_{\alpha}^{(n)},n=1,2,\ldots,6\}$. For each state $|n\rangle$, $\{Z_{\alpha}^{(n)},\alpha=1,2,\ldots,6\}$ (in the same color) signal the relative importance of $c\bar{c}$ and $D\bar{D}^{*}$ components.
To be specific, we determine the finite volume energies based on the correlation matrix of $c\bar{c}$ and $D\bar{D}^{*}$ operators, which compose an operator set $\tilde{\mathcal{S}}=\{\mathcal{O}_{\alpha}|\alpha=1,\cdots,6\}=\{\mathcal{O}_{c\bar{c}}^{r=0},\mathcal{O}_{c\bar{c}}^{r=1},\mathcal{O}_{c\bar{c}}^{r=2},\mathcal{O}_{D\bar{D}}^{\mathbf{q}=0,\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=0,\gamma_{4}\gamma_{5}},\mathcal{O}_{D\bar{D}}^{{q}=1}\}$. By solving the GEVP to the corresponding correlation matrix $C_{\alpha\beta}(t)$ similarly to Eq. (SI9), we obtain six eigenvectors $\{v_{\alpha}^{(n)},n=1,2,\ldots,6\}$ and six optimized correlation functions $\{C^{(n)}(t)=v_{\alpha}^{(n)}v_{\beta}^{(n)}C_{\alpha\beta}(t),n=1,2,\ldots,6\}$. It is expected that $C^{(n)}(t)$ is contributed most from the $n$-th eigenstate with the eigen-energy $E_{n}$ of the lattice Hamiltonian. The physical status of these states can be explored qualitatively by the coupling factor $Z_{\alpha}^{(n)}$ of $\mathcal{O}_{\alpha}$ to the $n$-th state, which is extracted through the expression
$Z_{\alpha}^{(n)}\equiv|\langle 0|\mathcal{O}_{\alpha}|n\rangle|\approx$ $\frac{\sqrt{2E_{n}}C_{\alpha\beta}(t^{*})v_{\beta}^{(n)}}{\sqrt{v_{\beta}^{(n)}v_{\gamma}^{(n)}C_{\beta\gamma}(t^{*})e^{-E_{n}t^{*}}}}$ (SI10)
at $t^{*}$ where the effective mass plateaus are almost reached. The relative coupling factors $Z_{\alpha}^{(n)}$ (normalized by the maximal value of $Z_{\alpha}^{(n)}$ for each operator $\mathcal{O}_{\alpha}$) are shown in Fig. SI3. Obviously, the $c\bar{c}$ operators $\mathcal{O}_{1,2,3}$ couple most to the $E_{1}$ state and also have substantial overlaps to the $E_{2}$ and $E_{3}$ states. In contrast, $\mathcal{O}_{4,5}$ ($D\bar{D}^{*}$ operators with $q=0$) couple mainly to $E_{2}$, and also overlap substantially to $E_{1}$ and $E_{3}$ states. $\mathcal{O}_{4}$ also has a large overlap to $E_{6}$. $\mathcal{O}_{6}$ ($\mathcal{O}_{D\bar{D}^{*}}(q=1)$ operator) couples predominantly to $E_{4}$ and has a sizable coupling to the $E_{3}$ state. Based on these information, we make the following raw identifiction:
  • The $E_{1}$ state of an energy around $3.5$ GeV is unambiguously assigned to be the conventional $\chi_{c1}(1P)$ state.
  • The $E_{6}$ state is coupled mainly by the $D\bar{D}^{*}(q=0)$ operator, but has a very high energy around 4.6-4.7 GeV (see the bottom rows of Tables SI1-SI4). According to the experimental evidence that the mass splitting of $1S-2S$ mesons are usually around 600 MeV, the $E_{6}$ state is very likely composed of a ground $D^{(*)}$ state and an excited $\bar{D}^{(*)}$ state with relative zero momentum.
  • The values of $E_{5}$ have fairly large errors and are insensitive to the inclusion or exclusion of $D\bar{D}^{*}$ operator. On the other hand, $E_{5}$ is relatively far from the energy region relevant to $X(3872)$, so we do not pay much attention to it.
  • The $E_{2},E_{3},E_{4}$ states are very relevant to $D\bar{D}^{*}$ scattering and $X(3872)$, therefore it is very crucial for them to be determined as precisely as possible.
Figure SI4: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods (Fit A, Fit B or Fit C in Eq. I) at $m_{\pi}=250$ MeV. The $x$-axis $t_{\mathrm{min}}$ denotes the different lower bounds of the fit window $[t_{\mathrm{min}},t_{\mathrm{max}}]$ (the upper bound is fixed to be $t_{\mathrm{max}}=35$). Top left: the fitting results of $E_{1}$ at different $t_{\mathrm{min}}$. The filled point indicates our final result. Top right: $E_{2}$ results. The grey band illustrates the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=0}$ (the $D\bar{D}^{*}$ threshold). The blue points show the two-exponential fit (Fit A) results of $E_{2}$, while the red points are the results $E_{2}=E_{D\bar{D}^{*}}^{q=0}+\Delta_{2}(0)$ from the ratio method (Fit B). The results of fit A and Fit B are consistent with each other for $t_{\mathrm{min}}/a_{t}\in[15,28]$ but the later have quite smaller statistical errors. So we take Fit B result at $t_{\mathrm{min}}/a_{t}=22$ (indicated by the filled point) as our final result for $E_{2}$ (the red band). Bottom left: $E_{3}$ results. The figure legends are similar to the case of $E_{2}$. Although consistency of the results of Fit A and Fit B, it is seen that the Fit B results has much smaller statistical errors and show little dependence on $t_{\mathrm{min}}$. So again we take Fit B result at $t_{\mathrm{min}}/a_{t}=17$ as our final result for $E_{2}$ (the red band). Bottom right: $E_{4}$ results: The grey band illustrates the non-interacting $D\bar{D}^{*}$ energy $E_{D\bar{D}^{*}}^{q=1}$. The blue points illustrate the Fit A result, while the green points are the results from Fit C (the ratio method) with $E_{4}=E_{D\bar{D}^{*}}^{q=0}+\Delta_{4}(1)$. The results of Fit C are consistent with the two-exponential fit (Fit A) results but have much smaller statistical errors. So we take the Fit C results at $t_{\mathrm{min}}/a_{t}=20$ as our final results (the green band).
Figure SI5: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods at $m_{\pi}=307$ MeV. The layout is similar to Fig. SI4. The filled point in each panel indicates the $t_{\mathrm{min}}/a_{t}$ where our final result is taken.
Figure SI6: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods at $m_{\pi}=362$ MeV. The layout is similar to Fig. SI4. The filled point in each panel indicates the $t_{\mathrm{min}}/a_{t}$ where our final result is taken.
Figure SI7: The self-consistency check of the fitted $E_{1,2,3,4}$ through different methods at $m_{\pi}=417$ MeV. The layout is similar to Fig. SI4. The filled point in each panel indicates the $t_{\mathrm{min}}/a_{t}$ where our final result is taken.
S3. Sysmatic uncertainties from anisotropic lattice
To facilitate a clearer comprehension with physical values in this study, the fixed $a_{t}^{-1}$ mean values are adopted to present the finite volume energies, notwithstanding the uncertainties associated with the lattice space $a_{s}$ and the anisotropy parameter $\xi\equiv a_{s}/a_{t}$. Besides, a sysmatic uncertainty on anisotropic parameter $\xi$ should be taken into account in finite volume scattering analysis. Generally, the finite volume scattering momenta are determined by solving the dispersion relation from prescribed energy levels,
$E_{n}=$ $\sqrt{m_{D}^{2}+\vec{p}_{n}^{2}}+\sqrt{m_{D^{*}}^{2}+\vec{p}_{n}^{2}},$ (SI11)
$\vec{p}_{n}^{2}=$ $\frac{1}{4}E_{n}^{2}-\frac{1}{2}(m_{D}^{2}+m_{D^{*}}^{2})+\frac{(m_{D}^{2}-m_{D^{*}}^{2})^{2}}{4E_{n}^{2}},$
So we use several fitting method to extract the values of $E_{2,3,4}$ and compare the results to check the self-consistency:
$\mathrm{Fit~{}A}:$ $C^{(n)}(t)=(A_{1}e^{-E_{n}t}+A_{2}e^{-E^{\prime}t})+(t\to(T-t)),$
$\mathrm{Fit~{}B}:$ $R^{(n)}(t,q=0)=\frac{C^{(n)}(t)}{C_{D}(t,q=0)C_{D*}(t,q=0)}\approx Ae^{-\Delta_{n}(0)t},$
$\mathrm{Fit~{}C}:$ $R^{(n)}(t.q=1)=\frac{C^{(n)}(t)}{C_{D}(t,q=1)C_{D*}(t,q=1}\approx Ae^{-\Delta_{n}(1)t}.$ (SI12)
Fit A is a two-state fitting. Fit B involves fitting the two-point corrector ratio $R(t)$ that $C(t)$ over the non-interacting $C_{D}(t,q=0)$ and $C_{D*}(t,q=0)$ using a single exponential function and then adding its non-interacting energy levels $E_{n}=E_{D\bar{D}^{*}}^{q}+\Delta_{n}(q)$ through the jackknife analysis. And Fit C is similar to Fit B but accounts for the non-interacting $C_{D}(t,q=1)$ and $C_{D*}(t,q=1)$. The comparisons are illustrated in Figs. SI4-SI7 for the four $m_{\pi}$’s. In each figure, the top-left panel shows the fitted results of $E_{1}$ at different $t_{\mathrm{min}}$ of the fit window $[t_{\mathrm{min}},t_{\mathrm{max}}]$. The maximum timeslice range $t_{\mathrm{max}}$ is set to 40 across all fits, as this is in the saturation range of these observed energy levels.
  • In the case of energy level $E_{1}$, it is far below the $D\bar{D}^{*}$ threshold and the simple two-states fitting perform well enough, while the ratio-exponential fits are unflavored due to the substantial effects of excited states artifacts.
  • The energy level $E_{2}$ is slightly below $D\bar{D}^{*}$ threshold about 20 MeV. The results form Fit A and Fit B reach commendable consistency.
  • For energy level $E_{3}$, the results form Fit B show greaer precision compared to Fit A (two-state fit). Consequently, the results from Fit B are adopted in our determination.
  • In the case of energy level $E_{4}$, Fit B (at relative momentum $q=0$) gives unstable results, as observed the energy level plateau exhibits a gradual slope, which suggesting the potential contamination of excited states . Conversely, Fit C is observed to gibe a more robust and dependable results for the energy plateau.
Practically, the dimensionless scattering momenta are derived from the relation $q^{2}=(\frac{2\pi}{\hat{L}\xi})^{2}(a_{t}\vec{p})^{2}$, where $(a_{t}\vec{p})^{2}$ expressed in lattice unit. In such a relation, it is necessary to take into account the uncertainty of the anisotropy parameter caused by solving the dimensionless scattering momenta. Since this study focuses on the $D\bar{D}^{*}$ channel, We choose to employ the anisotropy parameter measured from $D$ and $D^{*}$ meson. Given the slight deviation observed in the dispersion relations of the $D$ and $D^{*}$ mesons, as documented in Table SI5, we incorporate this discrepancy as a source of systematic uncertainty. The determined values $\xi_{\mathrm{sys.}}$ ensure that the one-sigma uncertainty range encompasses these systematic variations.
Table SI5: Anisotropy parameter $\xi$ and dispersion relation of various ensembles. The derivation of $\xi_{D}$ and $\xi_{D^{*}}$ is considered to be a sysmatic uncertainty in Eq. (SI11), noted as $\xi_{\mathrm{sys.}}$.
ens. $m_{\pi}/\mathrm{MeV}$ $a_{t}^{-1}/\mathrm{GeV}$ $\xi_{J/\psi}$ $\xi_{D}$ $\xi_{D^{*}}$ $\xi_{\mathrm{sys.}}$
M245 250(3) $7.276$ 5.061 (10) 5.070 (22) 5.126 (54) 5.11(7)
M305 307(2) $7.187$ 5.083 (10) 5.025 (16) 5.068 (30) 5.05(5)
M360 362(1) $7.187$ 4.989 (10) 4.974 (14) 4.990 (31) 4.99(3)
M415 417(1) $7.219$ 5.021 (11) 5.031 (16) 5.091 (39) 5.07(6)